# Brain functions emerge as thermal equilibrium states of the connectome

### v2

@@ Abstract
A fundamental paradigm in neuroscience is that cognitive functions—such as perception, learning, memory, and locomotion—are shaped by the brain’s structural organization. However, the theoretical principles explaining how this physical architecture governs its function remain elusive. Here, we propose an algebraic quantum mechanics (AQM) framework in which the functional states of a structural connectome emerge as thermal equilibrium states of an algebraic quantum system defined on the underlying directed multigraph. These equilibrium states, derived from the Kubo-Martin-Schwinger (KMS) states formalism, capture the contribution of each neuron to the overall information flow. We apply this framework to the connectome of the nematode Caenorhabditis elegans, providing a detailed description of the KMS states, exploring their functional implications, and predicting functional networks based on anatomical connectivity. Ultimately, our approach reveals functional circuits predicted by the topology of the connectome and illuminates on the mechanisms linking structure to function.
@@
The concept of KMS states on complex networks, originating from graph C*-algebras [14, 49] and the algebraic extension of quantum mechanics [24, 27], has been shown in [42] to provide statistical descriptors of the dynamic patterns of neuronal interactions. We investigate the functional implications of these descriptors and extract two fundamental features that characterize the functional organization of the C. elegans nervous system: (i) the pure functional connectome which maps the entire atlas of functional connections driven by the anatomical network, and (ii) an integration capacity (IC) index that measures the degree to which neurons can modulate multiple independent flows.
@@
Figure 1: Connectivity states of AFDR. Schematic visualizations of different types of connectivity states AFDR: structural connectivity (gray arrows), and the emittance networks (green and purple arrows) obtained from neural $\beta$-emittance profiles ($\operatorname{NEP_{\beta}}$) at different values of inverse temperature $\beta>\beta_{c}$, where $\beta_{c}$ is the critical inverse temperature. a: the structural connectivity of AFDR; the weights of the connections are the components of the structural connectivity vector $\operatorname{{\bf k}}^{AFDR}$. b: At $\beta_{s}=2.5\beta_{c}$, the emittance network coincides with the structural connectivity of AFDR. However, the connection to RMDVR (represented by the purple arrow) is not statistically significant ($p\sim 0.2$), and therefore not a pure functional connection (PFC). In other words, when the system is at this $\beta_{s}$–KMS state, AFDR can functionally communicates only with its direct neighbors at the same probability distribution as its structural connectivity state. c: At a relatively higher temperature $1/\beta$, long-range communications emerge (i.e. neurons that are not post-synaptic to AFDR, here positioned at the external concentric circle) while the neuron maintains neural emittances onto its direct neighbors (neurons positioned at the internal circle), albeit with lower intensity. d: At a very high temperature $1/\beta$ with $\beta$ close to the critical value, more longer range connections emerge, while some short range ones are either no longer present or not statistically significant. In both c and d, only the PFCs ($p<0.05$) are considered. Greens are sensory neurons, yellows are interneurons, and pinks are motor neurons.
Results
[…]
1. Neural emittance profiles. Given $\beta>0$, the emittance volume of $v$ at inverse temperature $\beta$ is defined as
@@ Results
which is an infinite series that encodes the volume of information flow downstream of $v$ when the system functions at inverse temperature $\beta$; i.e., when the functioning of each connection is affected by a factor of $e^{-\beta}$. We shall note that similar formulation was used for simple undirected graphs by Katz [32] to measure node status in sociometry, and Estrada and Hatano in [19] to study communicability in undirected complex networks. Observe that $Z^{\beta}_{v}\geq 1$, reflecting the fact that a node theoretically emits onto itself through the trivial path of length 00. In particular, if $v$ is not pre-synaptic to any neuron, its emittance volume is $1$. Furthermore, by an elementary result from functional analysis, there is a critical value $\beta_{c}=\log r=4.2958$, where $r$ is the spectral radius of $G$, such that $Z^{\beta}$ converges for $\beta>\beta_{c}$ and
$Z^{\beta}_{v}=\sum_{n=0}^{\infty}\sum_{u\in V}e^{-\beta n}\operatorname{{\textsf{A}}}^{n}_{uv}=\sum_{u\in V}(1-e^{-\beta}\operatorname{{\textsf{A}}})^{-1}_{uv},$ (2)
@@
We want emphasize that these vectors originate from the theory of KMS states in AQM [27] and graph C*-algebras [49]. Specifically, as we have shown in [42], information flow within the directed multigraph $G$ forms a system with infinitely many degrees of freedom that is best modeled by the graph C*-algebra of $G$. Furthermore, this graph C*–algebra carries a natural dynamics that turns it into an algebraic quantum system $\mathcal{O}_{G}$ whose KMS states at inverse temperature $\beta$—termed as $\beta$–KMS states—arise from statistical superpositions of the distributions $\operatorname{{\bf x}}^{v|\beta}$. Namely, a $\beta$–KMS state of the system $\mathcal{O}_{G}$ is associated to a convex combination
@@
where $(\operatorname{{\bf p}}_{v})_{v}$ is a distribution of relative quantities on neurons (see [42] for more details). It follows that the vectors $\operatorname{{\bf x}}^{v|\beta}$ can be interpreted as the pure $\beta$–KMS states or pure functional states of the connectome at inverse temperature $\beta$, while mixed $\beta$–KMS states are expressed via Eq. (4).
@@
One can illustrate pure functional states by representing the vector $\operatorname{{\bf x}}^{v|\beta}$ as a simple weighted directed star with source node $v$, where for $u\neq v$, $(v,u)$ is a directed edge of weight $\bar{\operatorname{{\bf x}}}^{v|\beta}_{u}=\operatorname{{\bf x}}^{v|\beta}_{u}/\sum_{w\neq v}\operatorname{{\bf x}}^{v|\beta}_{w}$ if $\operatorname{{\bf x}}^{v|\beta}_{u}>0$, thus obtaining a downstream connectivity network of the neuron consisting of potential functional connections, which we refer to as its emittance network.
For instance, Figs.1b–d represent the emittance networks of the thermosensory neuron $v={\rm AFDR}$ at different values of $\beta$, illustrating how they evolve as the (inverse) temperature varies. We see that increasing the temperature allows the emergence of long-range emittance connections which are facilitated by large amounts of parallel pathways of different lengths. Indeed, it is an immediate consequence of formula (3) that parallel paths and redundancies are critical to the emergence of strong neural emittances, and the lengths of pathways and the degree of redundancies that significantly contribute to their weights depend on temperature. More precisely, lower temperatures (i.e., higher values of $\beta$) favor parallel direct anatomical connections and short paths. Whereas, higher temperatures enhance the contributions of redundant long pathways of different lengths while lessening the importance of direct anatomical connections in the emittance weights. For example, AFDR has 13 synaptic connections and 52 anatomical paths of length 2 onto the interneuron AIYR, contributing greatly to the important weight of its neural emittance onto the latter at low temperature (Fig.1b). And at higher temperature values, while the neural emittance of AFRD $\longrightarrow$ AIYR decreases, other connections are ’established’ with neurons onto which AFDR can only be connected through thousands or millions of parallel anatomical paths of length $\geq 2$. For instance, there are 117 paths of length 2 and about 4 millions of length between 3 and 5 linking AFDR to RIAR, contributing to the neural emittance onto this neuron at $\beta=1.05\beta_{c}$ (Fig.1c). Similarly, the neural emittance to the ventral cord motor neuron VA11 that appears at $\beta=1.01\beta_{c}$ (Fig.1d) is largely due to the more than its 230 million incoming paths of length $\geq 5$ from AFDR.
@@
2. Structure-function divergence and structural temperature. How and to what degree do the emittance networks of a neuron ’differ’ from its structural connectivity? In order to address this question, let us consider the structural connectivity state of a neuron $v$ as the distribution $(\operatorname{{\bf k}}^{v}_{u})_{u}$ encoding the post-synaptic density of $v$; i.e., $\operatorname{{\bf k}}^{v}_{u}=\operatorname{{\textsf{A}}}_{uv}/k^{out}_{v}$ where $k^{out}_{v}=\sum_{w}\operatorname{{\textsf{A}}}_{wv}$ is the number of chemical and electrical outgoing connections from $v$. Then, using Uhlmann’s transition probability between quantum states [59, 29], we define the structure-function divergence (sfd) of a neuron $v$ as
@@
Specifically, $\operatorname{\bf sfd}$ measures how the two networks representing the $\operatorname{NEP_{\beta}}$ diverges from the anatomical wiring of the neuron in terms of both the number and intensity of their respective connections. Namely, it provides the ratio of non-overlapping links in both functional states. For instance, $\operatorname{\bf sfd}(AS08,\beta)=0.125$ for $\beta=1.7\beta_{c}$ means that, at this inverse temperature, the emittance network of the motor neuron AS08 deviates from anatomy by 12.5%.
Now $\operatorname{\bf sfd}(v,\beta)\sim 0$ if and only if the two distributions $\operatorname{{\bf k}}^{v}\sim\operatorname{{\bf x}}^{v|\beta}$, which would mean that when the system is in the pure functional state $\operatorname{{\bf x}}^{v|\beta}$, $v$ can send signals only to its direct neighbors at the same probabilities as in its structural connectivity state. We refer to the value of $\beta$ that minimizes $\max_{v}\operatorname{\bf sfd}(v,\beta)$ as the structural (inverse) temperature of the connectome, and will be denoted by $\beta_{s}$.
Since $\operatorname{\bf sfd}(v,\beta)\sim 0$ at $\beta=2.5\beta_{c}$ for each of the 280 neurons, we have $\beta_{s}\approx 10.7394$. In particular, going back to the example of AFDR, its neural $\beta_{s}$–emittance profile has approximately the same values as its structural connectivity state. For instance, the structural connection from AFDR to the motor neuron RMDVR has weight $\operatorname{{\bf k}}^{\small AFDR}_{\small RMDVR}=0.037037$, and its neural emittance onto RMDVR at $\beta_{s}$ is $\operatorname{{\bf x}}^{{\small AFDR}|\beta_{s}}_{\small RMDVR}\approx 0.037032$ (see Table 1). Namely, the emittance network of AFDR at inverse temperature $\beta_{s}$ perfectly maps over its anatomical connectivity network.
Analyzing $\operatorname{\bf sfd}$ variation of all neurons with respect to inverse temperature, we found that the emittance networks of some neurons, such as the motor neurons DD02, AVAL/R, AS04, remain close to their structural connectivity over long intervals, whereas those of other neurons, such as AS08, the polymodal sensory neurons PVDL/R and FLPL/R, and the interneurons LUAL/R, diverge rapidly from structure after certain low temperature values. For example, at $\beta=7.30$, AVAL and AVAR emittance networks deviate from structure by only 1.3%, while PVDL/R deviate by $\sim 9\%$ and AS08 deviates by 12.5% as mentioned earlier.
3. Functional temperatures. On the other hand, since every $\beta>\beta_{c}$ generates emittance networks and at values close to $\beta_{s}$ these networks do not considerably differ from anatomy, which $\beta$ values should we consider functionally relevant? To address this question, we examine the mean receptance at inverse temperature $\beta$; that is, the average probability for a neuron to receive signals from the other neurons, which is given by
@@
We then consider $\beta$ as a functional inverse temperature if the mean receptance is $>0.5$, meaning that at such a $\beta$, neurons receives more signals from the other neurons that from themselves through self-emittance. By computation we found that the mean receptance is 0.50.5~{}0.50.5 at $\beta_{\rm o}=1.07\beta_{c}$ and $>0.5$ on the interval $I_{f}=(\beta_{c},\beta_{\rm o})$. For instance, Fig. 1c and Fig. 1d represent the emittance networks of AFDR at two different functional inverse temperatures ($\beta=4.51$ and $\beta=4.34$, respectively).
[…]
$v$ $u$ $\operatorname{{\bf k}}^{v}_{u}$ $\operatorname{{\bf x}}^{v|\beta_{s}}_{u}$ p-value
AFDR ADFR 0.074074 0.074079 0.0006
ASHR 0.074074 0.074070 0.0002
URBR 0.037037 0.037031 0.0272
AIYR 0.481481 0.481441 0.0
ASEL 0.074074 0.074086 0.0
AFDL 0.074074 0.074073 0.0002
ASER 0.074074 0.074087 0.0004
AIBR 0.037037 0.037061 0.0472
AWCR 0.037037 0.037040 0.0046
RMDVR 0.037037 0.037032 0.1806
RMDVR OLQVL 0.0625 0.062489 0.0002
SAADL 0.03125 0.031243 0.0066
SIBDR 0.03125 0.031244 0.0008
SMBDR 0.03125 0.031244 0.0094
SAAVR 0.03125 0.031247 0.0056
CEPDR 0.03125 0.031247 0.002
RMDVL 0.03125 0.031272 0.0258
IL1L 0.03125 0.031246 0.0016
RMDR 0.03125 0.031256 0.0114
OLQDR 0.03125 0.031248 0.0018
RIAR 0.03125 0.031259 0.0416
IL1DR 0.0625 0.062498 0.0
RMDDL 0.125 0.125003 0.0
RMDDR 0.03125 0.031255 0.0192
SMDVR 0.03125 0.031251 0.0116
SIAVL 0.09375 0.093734 0.0
SIBVR 0.03125 0.031250 0.0034
URAVL 0.09375 0.093732 0.0
AFDR 0.03125 0.031243 0.0056
Table 1: Neural emittances of AFDR and RMDVR at $\beta_{s}$.
[…]
Figure 2: Emittance network of RID. The pink arrows are connections that are not determined by the network topology, that is, the particular anatomical wiring of the neuron post-synaptic connections, therefore not PFCs. Green arrows are emittance connections that are topology-dependent, therefore PFCs. These connections coincide with the functional connections from RID to the neuron classes URX and ADL that have been identified by Randi et al. to be driven by extrasynaptic signaling.
5. Putative extrasynaptic signaling pathways. A biological implication of emittance connections that are not PFCs is that the involved neurons do not necessarily rely on the fixed global anatomy to functionally interact with its target neurons, and additional ’structural’ connections might be needed for their functional connectivity ot be predicted from anatomy. This is consistent with results from multiple studies that have demonstrated the presence of extrasynaptic signaling between neurons [5, 36, 8, 51, 50]. To test this interpretation, we investigated the PFCs of the motor neuron and interneuron RID which has recently been identified by Randi et al. [50] to generate functional connections that were not predicted from anatomy.
On the other hand, considering an extrasynaptic chemical signaling as an actual directed physical connection between neurons, we separately added to the graph one edge from RID to URXL and one edge from RID to ADLR. And by recomputing the $\operatorname{NEP_{\beta}}$ of RID within the updated directed multigraphs, we found that (see Supplementary Data 5 and 6) the new neural emittances from RID to URXL and ADLR are PFCs ($p=0.0082$ and $p=0.0086$, respectively).
Figure 3: Pure functional network. Schematic of the pure functional connectome $\mathcal{F}^{\beta}$ of the C. elegans at a functional inverse temperature value $\beta_{f}=4.41$. It represents only $11.68\%$ of the whole directed weighted network defined by the matrix $[\operatorname{{\bf x}}^{\bullet|\beta}]$ whose column vectors are the neural $\beta$-emittance profiles of all neurons. Neurons are positioned according to their spatial coordinates [54].

### v3

@@ Abstract
A fundamental idea in neuroscience is that cognitive functions—such as perception, learning, memory, and locomotion—are shaped and constrained by the brain’s structural organization. Despite significant progress in mapping and analyzing structural connectomes, the principles linking the brain’s physical architecture to its functional capabilities remain elusive. Here, we introduce an algebraic quantum model to bridge this theoretical gap, offering new insights into the relationship between the connectome and emergent brain functions, while connecting structural data to functional predictions. Using the well-mapped C. elegans anatomical and extrasynaptic connectomes, we demonstrate that brain functions, defined as functional networks of a neural system, emerge as thermal equilibrium states of an algebraic quantum system derived from the graph algebra of the underlying directed multigraph. Specifically, these equilibrium states, characterized by the Kubo-Martin-Schwinger (KMS) formalism, reveal how individual neurons contribute to functional network formation. Our model illuminates the structure-function relationship in neural circuits through two key features: (1) a functional connectome that delineates topologically driven neuronal interactions and (2) an Integration Capacity (IC) index that quantifies how effectively neurons coordinate and modulate diverse information flows. Together, these features provide a statistical and mechanistic account of information flow and reveal how the network topology of the connectome predicts cognition and complex behaviors.
I Introduction
@@
The concept of KMS states on complex networks, originating from graph C*-algebras [29, 30] and the algebraic extension of quantum mechanics [31, 32], has been shown in [28] to provide statistical descriptors of the dynamic patterns of neuronal interactions. We investigate the functional implications of these descriptors and extract two fundamental features that characterize the functional organization of the C. elegans nervous system: (i) the purely topological functional connectome (PTFC) which maps the entire atlas of functional connections driven by the anatomical network, and (ii) an integration capacity (IC) index that measures the degree to which neurons can modulate multiple independent flows.
@@
The remainder of the paper is structured as follows. Section II introduces the theoretical framework and key derivations; Section III presents the main results on the structure-function relationship in C. elegans; and Section IV details the experimental validations. We conclude with a discussion of functional implications and broader perspectives in Section V.
II Theory
II.1 The connectome as a directed multigraph
We model the structure of the adult hermaphrodite C. elegans somatic connectome as a directed graph $G=(V,E)$ with parallel edges and self-loops, where the node set $V$ represents $N=280$ individual neurons, and the edge set $E$ consists of $12071$ unique chemical and electrical synapses (gap junctions), constructed by merging the old dataset [38, 13, 26] with the newly revised one [14] (see Methods for details and Supplementary Data 1 [39] for the resulting edgelist). The source $s(e)$ and range $r(e)$ of an edge are respectively the pre- and post-synaptic neurons of the synapse. Since gap junctions allow bidirectional transmission between neurons, they are represented by reciprocal edges. Our main focus being on the functional interactions among neurons through anatomical pathways, we consider both chemical and electrical synapses within a unified wiring system, since both transmission modalities are known to maintain close functional interactions [40]. We let $\operatorname{{\textsf{A}}}$ be the adjacency matrix of $G$. Specifically, since $G$ is a directed muligraph, $\operatorname{{\textsf{A}}}$ is an integer-weighted non-symmetric matrix, where for two neurons $u$ and $v$, the entry $\operatorname{{\textsf{A}}}_{uv}$ counts the number of chemical and electrical connections from $v$ to $u$. We use bold letters $\operatorname{\bf e},\operatorname{\bf f}$, etc. for directed walks; i.e., sequences $e_{1}\cdots e_{n}$ (with possible recurrences) of edges such that $s(e_{i})=r(e_{i+1})$. The length $|\operatorname{\bf e}|$ of such a walk $\operatorname{\bf e}$ is the number $n$ of edges composing it, and we set $s(\operatorname{\bf e})=s(e_{n})$ and $r(\operatorname{\bf e})=r(e_{1})$ to be its source and range, respectively. $\operatorname{\bf e}$ is a cycle if $s(\operatorname{\bf e})=r(\operatorname{\bf e})$. Since $G$ contains directed cycles and self-loops , finite length walks form an infinite set that we denoted by $E^{*}$. Indeed, 44 neurons have synapses onto themselves (such connections are referred to as ’autapses’ in the literature [41, 42]).
II.2 Neural emittance
Neural emittance profiles from KMS states.
Given $\beta>0$, the emittance volume of $v$ at inverse temperature $\beta$ is defined as
@@ II.2 Neural emittance
which is an infinite series that encodes the volume of information flow downstream of $v$ when the system functions at inverse temperature $\beta$; i.e., when the functioning of each connection is affected by a factor of $e^{-\beta}$. We shall note that similar formulation was used for simple undirected graphs by Katz [43] to measure node status in sociometry, and Estrada and Hatano in [44] to study communicability in undirected complex networks. Observe that $Z^{\beta}_{v}\geq 1$, reflecting the fact that a node theoretically emits onto itself through the trivial path of length 0. In particular, if $v$ is not pre-synaptic to any neuron, its emittance volume is $1$. Furthermore, by an elementary result from functional analysis, there is a critical value $\beta_{c}=\log r$, where $r$ is the spectral radius of $G$, such that $Z^{\beta}$ converges for $\beta>\beta_{c}$ and
$Z^{\beta}_{v}=\sum_{n=0}^{\infty}\sum_{u\in V}e^{-\beta n}(\operatorname{{\textsf{A}}}^{n})_{uv}=\sum_{u\in V}(1-e^{-\beta}\operatorname{{\textsf{A}}})^{-1}_{uv},$ (2)
@@
We want to emphasize that these vectors originate from the AQM methods [32] and the novel framework we developed for flow pathways in directed networks [28]. Specifically, as we have shown in [28], information flow within the directed multigraph $G$ forms a system with infinitely many degrees of freedom that is best modeled by the graph C*-algebra of $G$ [30]. Furthermore, this graph $C^{*}$–algebra carries a natural dynamics that turns it into an algebraic quantum system $\mathcal{O}_{G}$ whose KMS states at inverse temperature $\beta$—termed as $\beta$–KMS states—arise from statistical superpositions of the distributions $\operatorname{{\bf x}}^{v|\beta}$. Namely, a $\beta$–KMS state of the system $\mathcal{O}_{G}$ is associated to a convex combination
@@
where $(\operatorname{{\bf p}}_{v})_{v}$ is a distribution of relative quantities on neurons such as topological measures (out- or in-degree distributions), gene expression level of neurons, etc. (see [28] for theoretical details). It follows that the $\operatorname{NEP_{\beta}}$ vectors $\operatorname{{\bf x}}^{v|\beta}$ can be interpreted as the pure $\beta$–KMS states of the connectome at inverse temperature $\beta$, while mixed $\beta$–KMS states are expressed via Eq. (4).
To get a physical intuition of the notions of ”pure” and ”mixed” $\beta$-KMS states, think of the pure state $\operatorname{{\bf x}}^{v|\beta}$ as representing a state of the system where all the information flow originates from $v$, and the value $\operatorname{{\bf x}}^{v|\beta}_{u}$ quantifies the probability that this flow reaches $u$. In contrast, a mixed state $\operatorname{{\bf x}}$ of Eq. (4) describe a statistical ensemble where where information flow may originates from any neuron $v$ with probabily $\operatorname{{\bf p}}_{v}$. In this case, the value $\operatorname{{\bf x}}_{u}$ corresponds to the expected information flow received by neuron $u$ in this state under the probability distribution $(\operatorname{{\bf p}}_{v})_{v\in V}$.
@@
One can illustrate these states by representing the vector $\operatorname{{\bf x}}^{v|\beta}$ as a simple weighted directed star with source node $v$, where for $u\neq v$, $(v,u)$ is a directed edge of weight $\bar{\operatorname{{\bf x}}}^{v|\beta}_{u}=\operatorname{{\bf x}}^{v|\beta}_{u}/\sum_{w\neq v}\operatorname{{\bf x}}^{v|\beta}_{w}$ if $\operatorname{{\bf x}}^{v|\beta}_{u}>0$, thus obtaining a downstream connectivity network of the neuron consisting of potential functional connections, which we refer to as its emittance network.
Structure-function divergence.
@@
How and to what degree do the emittance networks of a neuron ’differ’ from its structural connectivity? In order to address this question, let us consider the structural connectivity state of a neuron $v$ as the distribution $(\operatorname{{\bf k}}^{v}_{u})_{u}$ encoding the post-synaptic density of $v$; i.e., $\operatorname{{\bf k}}^{v}_{u}=\operatorname{{\textsf{A}}}_{uv}/k^{out}_{v}$ where $k^{out}_{v}=\sum_{w}\operatorname{{\textsf{A}}}_{wv}$ is the number of chemical and electrical outgoing connections from $v$. Then, using Uhlmann’s transition probability between quantum states [46, 47], we define the structure-function divergence (sfd) of a neuron $v$ as
@@
Specifically, $\operatorname{\bf sfd}$ measures how the two networks representing the $\operatorname{NEP_{\beta}}$ diverges from the anatomical wiring of the neuron in terms of both the number and intensity of their respective connections.
Now $\operatorname{\bf sfd}(v,\beta)\sim 0$ if and only if the two distributions $\operatorname{{\bf k}}^{v}\sim\operatorname{{\bf x}}^{v|\beta}$, which would mean that when the system is in the functional state $\operatorname{{\bf x}}^{v|\beta}$, $v$ can send signals only to its direct neighbors at the same probabilities as in its structural connectivity state. We refer to the value of $\beta$ that minimizes $\max_{v}\operatorname{\bf sfd}(v,\beta)$ as the structural (inverse) temperature of the connectome, and will be denoted by $\beta_{s}$.
Functional temperatures.
On the other hand, since every $\beta>\beta_{c}$ generates emittance networks and at values close to $\beta_{s}$ these networks do not considerably differ from anatomy, which $\beta$ values should we consider functionally relevant? To address this question, we examine the mean receptance at inverse temperature $\beta$, defined as the average probability for a neuron to receive signals from the other neurons, and we compare it to self-emittance.
Specifically, for a fixed neuron $u$, consider the mixed $\beta$–KMS state $\operatorname{{\bf x}}(u,\beta)$ given by Eq. (4) where $\operatorname{{\bf p}}=(\operatorname{{\bf p}}_{v})$ is the uniform distribution on the neuron subset $\{v\in V;\ v\neq u\}$; namely, $\operatorname{{\bf p}}_{v}=1/(N-1)$ for $v\neq u$, and $\operatorname{{\bf p}}_{u}=0$. Then, $\operatorname{{\bf x}}(u,\beta)=\frac{1}{N-1}\sum_{v\neq u}\operatorname{{\bf x}}^{v|\beta}$, and the expected flow received by $u$ in this state is the value $\operatorname{{\bf x}}(u,\beta)_{u}=\frac{1}{N-1}\sum_{v\neq u}\operatorname{{\bf x}}^{v|\beta}_{u}$, which precisely measures the average information flow received by $u$ from all other neurons. Now, the mean receptance is obtained by taking the average $\langle\operatorname{{\bf x}}(u,\beta)_{u}\rangle=\frac{1}{N}\sum_{u}\operatorname{{\bf x}}(u,\beta)$, which is given by
@@
We then consider $\beta$ as a functional inverse temperature if the mean receptance is >0.5>0.5, meaning that at such a $\beta$, neurons receives more signals from the other neurons than from themselves through self-emittance.
II.3 Dependence on network topology
It is important to note that emittance networks result from global measures of macroscopic systems consisting of (possibly infinitely) many pathways of various depths, but they do not provide precise information about the extent to which particular synaptic circuitries are relevant to the emergence of these functional networks. More precisely, irrespective of its value, the weight of a connection in the emittance network does not, by itself, allow to know whether the occurrence of this connection is determined by the particular global topology of the anatomical network; i.e., the physical wiring of the connectome. For example, suppose that at a certain value of $\beta$ the neural emittance of $v$ onto $u$ coincides with an existing direct anatomical link. How can we know whether this particular physical wiring from $v$ to $u$ is ’essential’ for a this emittance to emerge at that value of $\beta$? And more generally, how dependent are the $\operatorname{NEP_{\beta}}$ on the network topology? To address these questions, we systematically compared the weights in the emittance networks of a given neuron to those that one would get if the network topology was randomly perturbed. Specifically, we measured the likelihood for a neural emittance obtained at a fixed value of $\beta$ to emerge with equal or larger weight at the same inverse temperature if the synaptic network was randomly rewired (see Methods). Based on their statistical significance, one can then distinguish the emittance connections that are determined by the network topology ($p<.05$) from those that are probably not. In a sense, the former represent the potential functional connections predicted by the network structure of the connectome; we then refer to them as purely topological functional connections (PTFC) of the neuron.
II.4 Purely topological functional connectomes from mixed states.